OpenStudy (anonymous):

help pls with 2 problems!!

OpenStudy (anonymous):

Select the system of equations that corresponds to the given graph.

OpenStudy (anonymous):

A)-2x - y = 4 2x + y = -2 B) -8x + 4y = -16 -4x + 2y = -8 C) x - 2y = 6 -2x + y = 4 D) 2x + 3y = 6 3x + 2y = 4

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

is it C?? thats what i got

OpenStudy (wolfe8):

I would put in x=0 to all eqations and see which pair gives y=2

OpenStudy (anonymous):

Awnser is D i did my work and the first one y equals 4 and the second one y equals -4 so its D @wolfe8

OpenStudy (wolfe8):

What?? But you want y=2 when x=0! See in the graph? The lines intercept at y=2 when x=0

OpenStudy (anonymous):

i dont know im confused i thought it would be D

OpenStudy (wolfe8):

Let me eat first

OpenStudy (anonymous):

ok

OpenStudy (wolfe8):

Ok I finished eating. So, take a look at the graph. Notice an important part of it - where both lines intersect each other at (0,2). So, since you are given choices, you can substitute x=0 into your options and see which pair both gives y=2.

OpenStudy (anonymous):

so its y = 2

OpenStudy (anonymous):

its C

OpenStudy (wolfe8):

Both of the equations have to give the same result. In C one of them gives y=-3

OpenStudy (anonymous):

oh so it has to be a or b @wolfe8

OpenStudy (wolfe8):

Are you just guessing? Did you even try what I told you to?

OpenStudy (anonymous):

yah plug in the numbers but now im guessing cause im getting confused my awnser was C but u said that wrongs

OpenStudy (wolfe8):

Replace x=0 into every single equation like this: -2(0)-y=4 2(0)+y=-2 And see what you get for y. The pair that gives you y=2 both is the answer.